OpenStudy (anonymous):

The circle with center O, the measure of minor arc JK is 124 degrees. What is the degree measure of angle OEK if JE is a diameter?

OpenStudy (anonymous):



OpenStudy (anonymous):

<JKE is right angle. Angle in a semi circle is 90 . JK and KE are radii , so equal in measure, <KJO and <OEK are equal

OpenStudy (anonymous):

I'm sorry, but what IS the measure? :(

OpenStudy (anonymous):

Angle JOK is 124. So angle KOE = 180 - 124 = 56. Triangle KOE is isosceles since KO=OE. Then angle OEK = (180 - 56) / 2 = 124 / 2 = 62 degrees.

OpenStudy (anonymous):

angle OEK is angle inscribed by arc JK which is half of measure of arc JK so angle OEK = 62 degrees